Curriculum vitæ — Jan KYBIC

July 8, 2008

Center for Machine Perception Department of Cybernetics Faculty of Electrical Engineering Czech Technical University Technická 2 Praha 6, 166 27 Czech Republic +420 2 2435 5721 (work) +420 2 2435 7385 (fax)

                                   kybic@fel.cvut.cz
                                   http://cmp.felk.cvut.cz/~kybic

Born on 18^{th} January 1974 in Prague, Czech Republic. Male.

Work experience

since 2003 Senior researcher at CMP, FEE, Czech Technical University, Prague.

2001–2003 Post-doc position at INRIA, France, working on inverse problems in MEG and EEG.

1992–1998 several part-time and summer jobs in software development and system administration.

Education

1998–2001 PhD in biomedical image processing at Ecole Polytechnique Fédérale de Lausanne (EPFL), Switzerland, with thesis Biomedical Image Registration by Elastic Warping.

1994–1998 Master (Engineer) in control engineering from Czech Technical University (CTU), with honors, with thesis Kalman Filtering and Speech Enhancement

Research interests

Signal and image processing for biomedical applications, image registration, segmentation, splines and wavelets, speech processing and enhancement, computer vision, numerical methods, inverse problems, algorithm theory, parallel algorithms, control theory.

Language skills

Mother tongue Czech, fluent English, French, and Spanish, conversational knowledge of German and Italian, intermediate Russian, basic Portuguese.

Professional activities

Associated Editor for IEEE Transactions on Medical Imaging. Article reviews for IEEE Transactions on Medical Imaging, IEEE Transactions on Image Processing, IEEE Transactions on Biomedical Engineering, Medical Image Analysis and various other journals and conferences. Member of IEEE, Czech TEX Users Group, Czech Linux Users Group. Organized the CVAMIA workshop at the ECCV conference in 2004, Prague; Local ACM Programming Contest at CTU in 1996–1999. Reviewer of PhD and Master theses. Collaborated on six projects funded by Czech grant agencies and in one European training network.

Pedagogical activities

Teaching a Digital Image Processing and Medical Imaging courses. Lectures about medical and microscopy image processing. Two PhD students and seven Master students have successfully defended; currently advising four and co-advising one PhD student.

Publications

Journal articles:

[1] Martin Barva, Marián Uherčík, Jean-Martial Mari, Jan Kybic, Jean-René Duhamel, Hervé Liebgott, Václav Hlaváč, and Christian Cachard, “Parallel integral projection transform for straight electrode localization in 3-d ultrasound images,” IEEE Transactions on Ultrasonics, Ferroelectrics, and Frequency Control (UFFC), vol. 55, no. 7, pp. 1559–1569, July 2008.

[2] Elisabeth Brusseau, Jan Kybic, Jean-François Déprez, and Olivier Basset, “2-D locally regularized tissue strain estimation from radio-frequency ultrasound images: Theoretical developments and results on experimental data,” IEEE Transactions on Medical Imaging, vol. 27, no. 2, pp. 145–160, Feb. 2008.

[3] Maureen Clerc and Jan Kybic, “Cortical mapping by Laplace-Cauchy transmission using a boundary element method.,” Journal on Inverse Problems, vol. 23, pp. 2589–2601, Nov. 2007.

[4] Jan Petr, Jan Kybic, Michael Bock, Sven Müller, and Václav Hlaváč, “Parallel image reconstruction using B-spline approximation (PROBER),” Magnetic Resonance in Medicine, vol. 58, no. 9, pp. 582–591, September 2007.

[5] Jan Kybic, Maureen Clerc, Olivier Faugeras, Renaud Keriven, and Théo Papadopoulo, “Generalized head models for MEG/EEG: BEM beyond nested volumes,” Physics in Medicine and Biology, vol. 51, no. 5, pp. 1333–1346, Feb. 2006, doi:10.1088/0031-9155/51/5/021.

[6] Jan Kybic, Maureen Clerc, Olivier Faugeras, Renaud Keriven, and Théo Papadopoulo, “Fast multipole acceleration of the MEG/EEG boundary element method,” Physics in Medicine and Biology, vol. 50, no. 19, pp. 4695–4710, October 2005.

[7] Olivier Faugeras, Geoffray Adde, Guillaume Charpiat, Christophe Chefd’Hotel, Maureen Clerc, Thomas Deneux, Rachide Deriche, Gerardo Hermosillo, Renaud Keriven, Pierre Kornprobst, Jan Kybic, Christophe Lenglet, Lucero Lopez-Perez, Théo Papadopoulo, Jean-Philippe Pons, Florent Segonne, Bertrand Thirion, David Tschumperle, Thierry Vieville, and Nicolas Wotawa, “Variational, geometric, and statistical methods for modeling brain anatomy and function,” Neuroimage, vol. 23S1, pp. S46–S55, 2004, Special issue: Mathematics in Brain Imaging.

[8] Jan Kybic, Maureen Clerc, Touffic Abboud, Olivier Faugeras, Renaud Keriven, and Théo Papadopoulo, “A common formalism for the integral formulations of the forward EEG problem,” IEEE Transactions on Medical Imaging, vol. 24, no. 1, pp. 12–28, Jan. 2005.

[9] María Jesús Ledesma-Carbayo, Jan Kybic, Manuel Desco, Andrés Santos, Michael Sühling, Patrick Hunziker, and Michael Unser, “Spatio-temporal non-rigid registration for ultrasound cardiac motion estimation,” IEEE Transactions on Medical Imaging, vol. 24, no. 9, pp. 1113–1126, Sept. 2005.

[10] Jan Kybic and Michael Unser, “Fast parametric elastic image registration,” IEEE Transactions on Image Processing, vol. 12, no. 11, pp. 1427–1442, November 2003.

[11] Jan Kybic, Thierry Blu, and Michael Unser, “Generalized sampling: A variational approach. Part I — Theory,” IEEE Transactions on Signal Processing, vol. 50, no. 8, pp. 1965–1976, August 2002.

[12] Jan Kybic, Thierry Blu, and Michael Unser, “Generalized sampling: A variational approach. Part II — Applications,” IEEE Transactions on Signal Processing, vol. 50, no. 8, pp. 1977–1985, August 2002.

[13] Jan Kybic, Philippe Thévenaz, Arto Nirkko, and Michael Unser, “Unwarping of unidirectionally distorted EPI images,” IEEE Transactions on Medical Imaging, vol. 19, no. 2, pp. 80–93, February 2000.

[14] Stanislav Tůma, Jiří Neuwirth, E. Čumlivská, J. Mališ, Jan Kybic, J. Šanda, M. Fricová-Poulová, T. Adla, and M. Polovinčák, “Nízkodávková technika spirální CT plic v diagnostice metastatických ložiskových nálezů u dětí a v dorostovém věku,” Česko-slovenská pediatrie, vol. 61, no. 4, pp. 179–185, Mar. 2006.

Books and book chapters:

[1] Tomáš Svoboda, Jan Kybic, and Hlaváč Václav, Image Processing, Analysis and Machine Vision — A MATLAB Companion, Thomson, Toronto, Canada, 1 edition, September 2007.

[2] Jan Kybic and Michael Unser, Handbook of Biomedical Image Analysis: Registration Models, vol. 3 of Topics in Biomedical Engineering, chapter Elastic Registration for Biomedical Applications, pp. 339–393, Kluwer Academic / Plenum Publishers, New York, USA, 2005.

[3] Milan Šonka, A. Ioannis Kakadiaris, and Jan Kybic, Eds., Computer Vision and Mathematical Methods in Medical and Biomedical Image Analysis (CVAMIA+MMBIA), number 3117 in Lecture Notes in Computer Science, Heidelberg, May 2004. Springer.

[4] Václav Hlaváč and Jan Kybic, Zdravotnická informatika, chapter Zpracování obrazu počítačem v medicíně, pp. 105–114, Karolinum, Univerzita Karlova v Praze, Ovocný trh 3–5, Praha 1, Sept. 2007, in Czech.

Conference articles:

[1] Jef Vandemeulebroucke, Patrick Clarysse, Jan Kybic, and David Sarrut, “Estimating respiratory motion from cone-beam projections,” in Proceedings of the First International Workshop on Pulmonary Image Analysis, Matthew Brown, Marleen de Bruijne, Bram van Ginneken, Atilla Kiraly, Jan-Martin Kuhnigk, Cristian Lorenz, Kensaku Mori, and Joseph Reinhardt, Eds., MorrisVille, USA, September 2008, pp. 83–92, Lulu.com.

[2] S. Tůma, J. Kybic, M. Dolejší, J. Polovinčák, J. Neuwirth, T. Adla, and E. Čumlivská, “Přesnost počítačem asistované CT diagnostiky plicních uzlů ve srovnání s klasickými postupy hodnocení,” in Čes.-slov. Pediatr., Jihlava, Czech Republic, Sept. 2008, vol. 63, pp. 52–31, Česká lékařská společnost J. E. Purkyně.

[3] T. Kazmar and J. Kybic, “Opacity quantification in cardiac angiogram sequences,” in BIOSIGNAL: Analysis of Biomedical Signals and Images, Jiří Jan, Jiří Kozumplík, and Ivo Provazník, Eds., Brno, Czech Republic, June 2008, p. 66, VUTIUM Press, CDROM proceedings.

[4] Marián Uherčík, Jan Kybic, Hervé Liebgott, and Christian Cachard, “Multiresolution parallel integral projection for fast localization of a straight electro de in 3d ultrasound images,” in Proceedings of 2008 IEEE International Symposium on Biomedical Imaging: From Nano to Macro, New York, US, May 2008, IEEE International Symposium on Biomedical Imaging (ISBI), pp. 33–36, IEEE Press.

[5] Jan Kybic, “Fast no ground truth image registration accuracy evaluation: Comparison of bootstrap and Hessian approaches,” in Proceedings of 2008 IEEE International Symposium on Biomedical Imaging: From Nano to Macro, New York, US, May 2008, pp. 792–795, IEEE Press.

[6] Martin Dolejší, Jan Kybic, Michal Polovinčák, and Stanislav Tůma, “Reducing false positive responses in lung nodule detector system by Asymmetric AdaBoost,” in Proceedings of 2008 IEEE International Symposium on Biomedical Imaging: From Nano to Macro, 3 Park Ave., 17th Floor, New York, NY 10016-5997, USA, May 2008, pp. 656–659, IEEE.

[7] Juan D. García-Arteaga and Jan Kybic, “Regional image similarity criteria based on the Kozachenko-Leonenko entropy estimator,” in CVPRW: Computer Vision and Pattern Recognition Workshops, 445 Hoes Lane, Piscataway, NJ, U.S.A., June 2008, pp. 1–8, IEEE.

[8] Jakub Krátký and Jan Kybic, “Three-dimensional segmentation of bones from CT and MRI using fast level sets,” in SPIE MI 2008: Proceedings of the SPIE Medical Imaging 2008 Conference, Joseph M. Reinhardt and Josien P. W. Pluim, Eds., Bellingham, Washington, USA, February 2008, SPIE, vol. 6914 of Medical Imaging 2008: Image Processing, p. 10, SPIE, CD-ROM.

[9] Jan Kybic, “High-dimensional entropy estimation for finite accuracy data: R-NN entropy estimator,” in IPMI2007: Information Processing in Medical Imaging, 20th International Conference, Nico Karssemeijer and Boudewijn Lelieveldt, Eds., Berlin, Heidelberg, Germany, July 2007, pp. 569–580, Springer.

[10] Jan Petr and Jan Kybic, “Continuous criterion for parallel MRI reconstruction using B-spline approximation (PROBER),” in Proceedings of SPIE, Josien P. W. Pluim and Joseph M. Reinhardt, Eds., Bellingham, Washington, USA, February 2007, SPIE, vol. 6512 of Medical Imaging 2007: Image processing, pp. 43–53, SPIE.

[11] Juan D. García-Arteaga and Jan Kybic, “Automatic landmark detection for cervical image registration validation,” in Proceedings of SPIE, Maryellen L. Giger and Nico Karssemeijer, Eds., Bellingham, Washington, USA, February 2007, SPIE, vol. 6514 of Medical Imaging 2007: Computer-Aided Diagnosis, p. 65142S, SPIE.

[12] Juan D. García-Arteaga, Jan Kybic, Jia Gu, and Wenjing Li, “Geometric and information constraints for automatic landmark selection in colposcopy sequences,” in VISAPP 2007: Proceedings of the Second International Conference on Computer Vision Theory and Applications, Alpesh Kumar Ranchordas, Helder Araújo, and Jordi Vitria, Eds., Setúbal,Portugal, March 2007, INSTICC, pp. 333–338, INSTICC-Institute for Systems and Technologies of Information, Control and Communication.

[13] Jan Kybic, Olivier Faugeras, Maureen Clerc, and Theo Papadopoulo, “Neural mass model parameter identification for MEG/EEG,” in Proceedings of SPIE: Medical Imaging 2007. Physiology, Function, and Structure from Medical Imaging, A. Manduca and P. Hu Xiaoping, Eds., Bellingham, Washington, USA, Feb. 2007, SPIE, vol. 6511, pp. 0F–1–0F–9, SPIE.

[14] Martin Dolejší and Jan Kybic, “Automatic two-step detection of pulmonary nodules,” in Proceedings of SPIE, Maryllen L. Giger and Nico Karssemeijer, Eds., Bellingham, Washington, USA, February 2007, vol. 6514 of Medical Imaging 2007: Computer-Aided Diagnosis, pp. 3j–1–3j–12, SPIE.

[15] S. Tůma, J. Neuwirth, M. Dolejší, J. Kybic, K. Daníčková, M. Polovinčák, J. Šanda, E. Čumlivská, and J. Mališ, “Možnost snižování dávky ionizujícího záření při vyšetření nádorových onemocnění plic u dětí a dorostu.,” in Čes.-slov. Pediatr., Praha, Czech Republic, June 2007, number 5, pp. 357–358, Česká lékařská společnost J. E. Purkyně.

[16] Daniel Smutek, Radim Šára, Štěpán Holinka, Jan Kybic, Ludvík Tesař, Jan Jiskra, and Pavel Maruna, “Ultrasound image of chronic thyroiditis and its relation to antithyroid antibodies,” in Proceedings of the 11th Congress of the World Federation for Ultrasound in Medicine and Biology, Amsterdam, Holland, May-June 2006, World Federation for Ultrasound in Medicine and Biology, vol. 32 of Ultrasound in Medicine and Biology, p. 120, Elsevier.

[17] Jan Kybic and Daniel Smutek, “Image registration accuracy estimation without ground truth using bootstrap,” in CVAMIA: Computer Vision Approaches to Medical Image Analysis, R. Beichel and M. Sonka, Eds. May 2006, number 3117 in Lecture Notes in Computer Science, pp. 61–72, Springer.

[18] Ignacio Arganda-Carreras, Carlos O. S. Sorzano, Roberto Marabini, Jose M. Carazo, Carlos Ortiz de Solorzano, and Jan Kybic, “Consistent and elastic registration of histological sections,” in CVAMIA: Computer Vision Approaches to Medical Image Analysis, R. Beichel and M. Sonka, Eds., Heidelberg, Germany, May 2006, number 3117 in Lecture Notes in Computer Science, pp. 85–95, Springer, http://dx.doi.org/10.1007/11889762 8.

[19] Jan Kybic, “Incremental updating of nearest neighbor-based highdimensional entropy estimation,” in ICASSP2006, Pierre Duhamel and Luc Vandendorpe, Eds., Toulouse, France, May 2006, pp. III–804, IEEE, DVD proceedings.

[20] Martin Dolejší and Jan Kybic, “Detection of pulmonary nodules in CT scans,” in BIOSIGNAL: Analysis of Biomedical Signals and Images, Jiří Jan, Jiří Kozumplík, and Ivo Provazník, Eds., Brno, Czech Republic, June 2006, pp. 251–253, VUTIUM Press.

[21] Jan Petr, Jan Kybic, Václav Hlaváč, Sven Müller, and Michael Bock, “Fast parallel MRI reconstruction using B-spline approximation (PROBER),” in Proceedings of SPIE, Michael J. Flynn and Jiang Hsieh, Eds., Bellingham, Washington, USA, February 2006, SPIE, vol. 6142 of Medical Imaging 2006: Physics of Medical Imaging, pp. 1251–1262, SPIE.

[22] Jan Petr, Jan Kybic, Sven Müller, Michael Bock, and Václav Hlaváč, “Parallel MRI reconstruction using B-spline approximation,” in ISMRM ’05: Proceedings of the 13th Scientific Meeting and Exhibition of International Society for Magnetic Resonance in Medicine, 2118 Milvia Street, Berkeley, USA, May 2005, ISMRM, vol. 13, p. 2421, International Society for Magnetic Resonance in Medicine.

[23] Dimitris Glotsos, S. Tsantis, Jan Kybic, A. Daskalakis, P. Ravazoula, I. Kalatzis, N. Dimitropoulos, D. Cavouras, and G. Nikiforidis, “Maxima chain edge representation for nuclei detection in microscopy images of thyroid nodules,” in EMBEC 2005, Jiří Hozman Peter Kneppo, Ed., Argentinska 38, 170 00, Prague, Czech Republic, November 2005, p. 0, International Federation for Medical and Biological Engineering.

[24] Martin Barva, Jan Kybic, Jean-Martial Mari, Christian Cachard, and Václac Hlaváč, “Localizing metal electrode from 3D ultrasound data using RANSAC and intensity priors,” in IFMBE Proceedings EMBEC’05, 3rd European Medical and Biological Engineering Conference, Jiří Hozman Peter Kneppo, Ed., Argentinska 38, 170 00, Prague, Czech Republic, November 2005, vol. 11, p. 6, International Federation for Medical and Biological Engineering, proceeding on CD only.

[25] María-Jesús Ledesma-Carbayo, Andrés Santos, Patricia Mahia, Miguel-Angel Garcia-Fernandez, Jan Kybic, Malpica Norberto, Perez-David E., and Desco Manuel, “Longitudinal and radial regional strain obtained from gray-scale conventional echocardiography,” in Journal of the American College of Cardiology, B. Lerman, Ed., Bethesda, USA, Mar. 2005, vol. 45, p. A255, Elsevier, supplement 1.

[26] Leila Muresan, Bettina Heise, Jan Kybic, and Erich Peter Klement, “Quantitative analysis of microarrays,” in Proceedings of ICIP, 445 Hoes Lane, Piscataway, USA, September 2005, pp. 1274–1277, IEEE.

[27] Jan Kybic and Daniel Smutek, “Computational elastography from standard ultrasound image sequences by global trust region optimization,” in Proceedings of IPMI, Lecture Notes in Computer Science, Milan Šonka and Gary Christensen, Eds., Heidelberg, Germany, July 2005, number 3565, pp. 299–310, Springer Verlag.

[28] Jan Kybic and Daniel Smutek, “Estimating elastic properties of tissues from standard 2d ultrasound images,” in Medical Imaging 2005: Ultrasonic Imaging and Signal Processing, William F. Walker and Stanislav Y. Emelianov, Eds., Bellingham, Washington, USA, February 2005, vol. 6 of Progress in Biomedical Optics and Imaging, pp. 184–195, SPIE.

[29] Maureen Clerc, Geoffray Adde, Jan Kybic, Théo Papadopoulo, and Jean-Michel Badier, “In vivo conductivity estimation with symmetric boundary elements,” in NFSI2005: 5th International Conference on Bioelectromagnetism and 5th International Symposium on Noninvasive Functional Source Imaging. May 2005, vol. 7, International Society for Bioelectromagnetism.

[30] Martin Barva, Jan Kybic, Jean-Martial Mari, Christian Cachard, and Václav Hlaváč, “Automatic localization of curvilinear object in 3D ultrasound images,” in Medical Imaging 2005: Ultrasonic Imaging and Signal Processing, William F. Walker and Stanislav Y. Emelianov, Eds., Bellingham, Washington, USA, February 2005, SPIE International Symposium on Medical Imaging, vol. 6 of Progress in Biomedical Optics and Imaging, pp. 455–462, SPIE.

[31] Maureen Clerc, Jean-Michel Badier, Geoffray Adde, Jan Kybic, and Théo Papadopoulo, “Boundary element formulation for electrical impedance tomography,” in CEMRACS 2004: Mathematics and applications to biology and medicine, E. Cancès and J.-F.

Gerbeau, Eds., 17, avenue du Hoggar, France, July–September 2004, vol. 14 of ESAIM: Proceedings, pp. 63–71, EDP Sciences, http://www.edpsciences.org/articles/proc/abs/2005/01/contents/contents.html.

[32] Jan Kybic, “High-dimensional mutual information estimation for image registration,” in ICIP’04: Proceedings of the 2004 IEEE International Conference on Image Processing, 445 Hoes Lane, Piscataway, NJ, U.S.A., October 2004, IEEE Computer Society.

[33] Jan Kybic, Maureen Clerc, Olivier Faugeras, Geoffray Adde, Renaud Keriven, and Théo Papadopoulo, “Accurate boundary element method for the electro- and magnetoencephalography forward problem,” in BIOSIG-NAL 2004: 17-th Biennial International Eurasip Conference Proceedings, Jiří Jan, Jiří Kozumplík, and Ivo Provazník, Eds., Brno, Czech Republic, June 2004, pp. 121–123, VUTIUM Press.

[34] Jan Petr, Václav Hlaváč, Yehudah L. Werner, and Jiří Moravec, “Tools for Image Analysis of Gecko Skin,” in BIOSIGNAL 2004: Analysis of Biomedical Signals and Images, Jiří Jan, Jiří Kozumplík, and Ivo Provazník, Eds., Brno, Czech Republic, June 2004, EURASIP, vol. 17, pp. 296–298, VU-TIUM Press.

[35] Martin Barva, Jean-Martial Mari, Jan Kybic, and Christian Cachard, “Radial Radon transform for electrode localization in biological tissue,” in BIOSIGNAL: Analysis of Biomedical Signals and Images, Jiří Jan, Jiří Kozumplík, and Ivo Provazník, Eds., Brno, Czech Republic, June 2004, EURASIP, pp. 299–301, VUTIUM Press.

[36] Martin Barva, Jan Kybic, Jean-Martial Mari, and Christian Cachard, “Radial Radon transform dedicated to micro-object localization from radio frequency ultrasound signal,” in IEEE UFFC: Proceedings of the IEEE International Frequency Control Symposium and Exposition, Marjorie Passini Yuhas, Ed., 445 Hoes Lane, Piscataway, NJ 08855-1331 USA, August 2004, IEEE, pp. 1836–1839, IEEE Operations Center.

[37] María Jesús Ledesma-Carbayo, Andrés Santos, Patricia Mahia, Miguel-Angel García Fernández, Jan Kybic, Norberto Malpica, E. Pérez David, and Manuel Desco, “Longitudinal and radial regional strain obtained from gray-scale conventional echocardiography,” in Abstracts of EUROECHO, the Eighth Annual Meeting of the European Association of Echocardiography, Dec. 2004, number 1001, p. S151.

[38] María Jesús Ledesma-Carbayo, Manuel Desco, Norberto Malpica, Jan Kybic, and Andrés Santos, “Non rigid registration for myocardial motion analysis,” in 4th European Symposium on Biomedical Engineering (CD-ROM), June 2004.

[39] María Jesús Ledesma-Carbayo, Andrés Santos, P. Mahía Casado, M. A. García Fernández, J. Kybic, E. Pérez David, M. Desco, and F. Sarnago Cebada, “Obtención de las componentes axial y longitudinal del strain miocárdico a través de ecocardiografía convencional en escala de grises,” in XX Congreso Nacional de la SEC. Rev. Esp. Cardiol., 2004, number 67 (Supl 2).

[40] Jan Kybic and Maureen Clerc, “Symmetric BEM and multiscale fast multipole method for the E/MEG problem,” in NFSI 2003: Proceedings of the 4th International Symposium on Noninvasive Functional Source Imaging Within the Human Heart and Brain, V. Pizzella and G. L. Romani, Eds., Berlin, Germany, Sept. 2003, pp. 122–124, Organ of the German Society for Biomedical Engineering in VDE and Swiss and Austrian Societies for Medical and Biomedical Engineering.

[41] Geoffray Adde, Maureen Clerc, Renaud Keriven, and Jan Kybic, “Regularization for the inverse EEG/MEG problem using the symmetric boundary element method,” in Proceedings of the 4th International Symposium on Noninvasive Functional Source Imaging (NFSI), V. Pizzella and G. L. Romani, Eds., Berlin, Germany, September 2003, pp. 287–289, Organ of the German Society for Biomedical Engineering in VDE.

[42] Geoffray Adde, Maureen Clerc, Olivier Faugeras, Renaud Keriven, Jan Kybic, and Théo Papadopoulo, “Symmetric BEM formulation for the M/EEG forward problem,” in Proceedings of IPMI, Lecture Notes in Computer Science, Alison Noble and Chris Taylor, Eds., Heidelberg, Germany, July 2003, pp. 524–535, Springer Verlag.

[43] Maureen Clerc, Alain Dervieux, Renaud Keriven, Olivier Faugeras, Jan Kybic, and Théo Papadopoulo, “Comparison of BEM and FEM methods for the E/MEG problem,” in Biomag 2002: Proceedings of the 13th International Conference on Biomagnetism, Hannes Nowak, Jens Haueisen, Frank Giessler, and Ralph Huonker, Eds., Germany, August 2002, pp. 688–690, VDE Verlag GmbH.

[44] Maureen Clerc, Renaud Keriven, Olivier Faugeras, Jan Kybic, and Théo Papadopoulo, “The fast multipole method for the direct E/MEG problem,” in Proceedings of the First 2002 IEEE International Symposium on Biomedical Imaging: Macro to Nano (ISBI’02), 445 Hoes Lane, Piscataway, NJ, U.S.A., July 2002, pp. 1023 – 1026, IEEE.

[45] María Jesús Ledesma-Carbayo, Jan Kybic, Michael Sühling, P. Hunziger, Manuel Desco, Andrés Santos, and Michael Unser, “Cardiac ultrasound motion detection by elastic registration exploiting temporal coherence,” in Proceedings of the First 2002 IEEE International Symposium on Biomedical Imaging: Macro to Nano (ISBI’02), 445 Hoes Lane, Piscataway, NJ, U.S.A., July 2002, pp. 585 – 588, IEEE.

[46] María Jesús Ledesma-Carbayo, J. Kybic, M. Desco, E. Pérez-David, M. Unser, and A. Santos, “Evaluación de la funcionalidad regional del ventrículo izquierdo mediante técnicas de registro elástico espaciotemporal,” in XX Congreso Anual de la Soc. Esp. Ing. Biomédica. November 2002, pp. 187–190, Univ. Zaragoza.

[47] Maureen Clerc, Olivier Faugeras, Renaud Keriven, Jan Kybic, and TPéo Papadopoulo, “A level set method for the inverse EEG/MEG problem,” in Human Brain Mapping – 8th International Conference on Functional Mapping of the Human Brain, June 2002, Available on CD-Rom in NeuroImage, Vol. 16, No. 2.

[48] Jan Kybic, Thierry Blu, and Michael Unser, “Generalized sampling: A variational approach,” in Proceedings of the Fourth International Conference on Sampling Theory and Applications (SampTA’01), May 2001, pp. 151– 154.

[49] Jan Kybic, Thierry Blu, and Michael Unser, “Variational approach to tomographic reconstruction,” in Proceedings of the SPIE International Symposium on Medical Imaging: Image Processing (MI’01), Part I, M. Sonka and K.M. Hanson, Eds., February 2001, vol. 4322 of Progress in Biomedical Optics and Imaging, pp. 30–39.

[50] María Jesús Ledesma-Carbayo, Jan Kybic, Manuel Desco, Andrés Santos, and Michael Unser, “Cardiac motion analysis from ultrasound sequences using non-rigid registration,” in Proceedings of MICCAI, Wiro J. Niessen and Max A. Viergever, Eds. October 2001, pp. 889–896, Springer-Verlag.

[51] Jan Kybic and Michael Unser, “Multidimensional elastic registration of images using splines,” in ICIP2000, Proceedings of International Conference of Image Processing, September 2000, vol. II, pp. 455–458.

[52] Jan Kybic, Philippe Thévenaz, and Michael Unser, “Compensation of unidirectional geometric distortion in EPI using spline warping,” in ICIP’99: Proceedings of the 1999 IEEE International Conference on Image Processing, 445 Hoes Lane, Piscataway, NJ, U.S.A., October 1999, vol. II, pp. 168–172, IEEE Computer Society.

[53] Jan Kybic, Philippe Thévenaz, and Michael Unser, “Multiresolution spline warping for EPI registration,” in Proceedings of the SPIE Conference on Mathematical Imaging: Wavelet Applications in Signal and Image Processing VII. July 1999, vol. 3813, pp. 571–579, SPIE.

[54] Pavel Sovka, Petr Pollák, and Jan Kybic, “Extended spectral subtraction,” in Proceedings of EUSIPCO’96, 1996, vol. 2, p. 963.

Supervised Master and PhD theses:

[1] Martin Barva, Localization of Surgical Instruments in 3D Ultrasound Images, PhD Thesis CTU–CMP–2007–12, Center for Machine Perception, K13133 FEE Czech Technical University, Prague, Czech Republic, June 2007.

[2] Jan Petr, Parallel Magnetic Resonance Imaging Reconstruction, Ph.D. thesis, Czech Technical University in Prague, Prague, Czech Republic, October 2007.

[3] Helena Jansová, “Reconstruction of 3D surface from colonoscopic video,” M.S. thesis, Center for Machine Perception, K13133 FEE Czech Technical University, Prague, Czech Republic, May 2008.

[4] Jiří Mazanec, “Program pro prohlížení a segmentaci 3D MRI dat,” M.S. thesis, Center for Machine Perception, K13133 FEE Czech Technical University, Prague, Czech Republic, May 2008.

[5] Tomáš Kazmar, “Opacity quantification in cardiac angiogram sequence,” M.S. thesis, Faculty of Mathematics and Physics, Charles University, Prague, Czech Republic, May 2008.

[6] Marek Jasanský, “Paralelizace řešení přímého problému EEG rekonstrukce metodou BEM,” MSc Thesis CTU–CMP–2007–11, Center for Machine Perception, K13133 FEE Czech Technical University, Prague, Czech Republic, June 2007.

[7] Jiří Svoboda, “Porovnání metod odhadování entropie pro registraci obrazů,” MSc Thesis CTU–CMP–2007–11, Center for Machine Perception, K13133 FEE Czech Technical University, Prague, Czech Republic, May 2007.

[8] Martin Dolejší, “Detection of pulmonary nodules from CT scans,” MSc Thesis CTU–CMP–2007–05, Center for Machine Perception, K13133 FEE Czech Technical University, Prague, Czech Republic, January 2007.